# Enhancing Efficiency and Exploration in Reinforcement Learning for LLMs

### v1

@@ Abstract
Mengqi Liao1,2, Xiangyu Xi2, Ruinian Chen2, Jia Leng2, Yangen Hu2, Ke Zeng2, Shuai Liu1, Huaiyu Wan 1Corresponding author, 1Department of Computer Science, Beijing Jiaotong University, 2Meituan, Correspondence: hywan@bjtu.edu.cn
@@ 2 Related Work
Exploration and Exploitation in Reinforcement Learning. Balancing exploration and exploitation is a central challenge in reinforcement learning. Common strategies include $\varepsilon$-greedy, Upper Confidence Bounds (UCB), and Boltzmann Exploration Sutton et al. (1998). Boltzmann Exploration selects actions based on a softmax probability distribution proportional to the exponential of the estimated values of actions, regulated by a temperature parameter $\tau$. Similarly, LLMs generate tokens using a softmax distribution. Asadi and Littman (2017) introduced the Mellowmax operator to enhance the stability of Softmax, while Kim and Konidaris (2019) applied meta-gradient reinforcement learning to dynamically adjust temperature parameter of Mellowmax for better exploration-exploitation trade-offs. Moreover, Entropy Regularization, a common technique in deep reinforcement learning, adds an entropy term to the optimization objective to encourage stochastic policies and broader exploration Williams (1992).
@@ 4.2 Temperature Scheduling to Promote Exploration
Formula (4) ensures that the scaling factor of entropy, after temperature adjustment, maintains an approximately linear relationship with $\alpha$, as illustrated in Figure 2. This indicates that entropy returns to the level of $H_{\text{init}}$ after the temperature adjustment. The detailed derivation of Formula (4) is provided in Appendix A. The temperature scheduler maintains the policy’s entropy consistently at a stable level, as shown in Figure 1, thereby effectively enhancing policy exploration.
@@
$H_{\text{anneal}}^{(t)}$ $=H_{\text{init}}\cdot\left[\eta+(1-\eta)\cdot\frac{1}{2}\left(1+\right.\right.$
@@ Appendix A Entropy Scaling through Temperature Adjustment in Softmax Distributions
where $\beta=1/\tau$ and $Z(\beta)=\sum_{j=1}^{N}e^{-\beta\Delta_{j}}$The Shannon entropy Shannon (1948) of this distribution is defined as:
@@
When the entropy is small, the probability corresponding to $z_{\text{max}}$ is given by $p_{\text{max}}=\frac{1}{Z(\beta)}\approx 1$, which implies that $Z(\beta)\approx 1$. Consequently, for the remaining $N-1$ states, the probabilities can be approximated as:

### v2

@@ Abstract
Mengqi Liao1,2, Xiangyu Xi2, Ruinian Chen2, Jia Leng2, Yangen Hu2, Ke Zeng2, Shuai Liu1, Huaiyu Wan 1,3***Corresponding author, 1School of Computer Science and Technology, Beijing Jiaotong University 2Meituan 3Beijing Key Laboratory of Traffic Data Mining and Embodied Intelligence *Correspondence: hywan@bjtu.edu.cn
@@ 2 Related Work
Exploration and Exploitation in Reinforcement Learning. Balancing exploration and exploitation is a central challenge in reinforcement learning. Common strategies include $\varepsilon$-greedy, Upper Confidence Bounds (UCB), and Boltzmann Exploration Sutton et al. (1998). Boltzmann Exploration selects actions based on a softmax probability distribution proportional to the exponential of the estimated values of actions, regulated by a temperature parameter $\tau$. Similarly, LLMs generate tokens using a softmax distribution. Asadi and Littman (2017) introduced the Mellowmax operator to enhance the stability of Softmax, while Kim and Konidaris (2019) applied meta-gradient reinforcement learning to dynamically adjust temperature parameter of Mellowmax for better exploration-exploitation trade-offs. Moreover, Entropy Regularization, a common technique in deep reinforcement learning, adds an entropy term to the optimization objective to encourage stochastic policies and broader exploration Williams (1992). Apart from these reinforcement learning-based methods, a self-improvement approach Zeng et al. (2024) similarly enhances training performance by maintaining the exploration capability of the LLM through adjustments among several discrete temperature levels.
@@ 4.2 Temperature Scheduling to Promote Exploration
Formula (4) ensures that the scaling factor of entropy, after temperature adjustment, maintains an approximately linear relationship with $\alpha$, as illustrated in Figure 2. This indicates that entropy returns to the level of $H_{\text{init}}$ after the temperature adjustment. The detailed derivation of Formula (4) is provided in Appendix A. The temperature scheduler maintains the policy’s entropy consistently at a stable level, as shown in Figure 1, thereby effectively enhancing policy exploration. Moreover, it is important to note that logits are also divided by the temperature during forward propagation after sampling, ensuring consistency between the LLM’s distribution during training and sampling.
@@
[…]
@@ Appendix A Entropy Scaling through Temperature Adjustment in Softmax Distributions
where $\beta=1/\tau$ and $Z(\beta)=\sum_{j=1}^{N}e^{-\beta\Delta_{j}}$ .The Shannon entropy Shannon (1948) of this distribution is defined as:
@@
When the entropy is small, the probability corresponding to $z_{\text{max}}$ is given by $p_{\text{max}}=\frac{1}{Z(\beta)}\approx 1$, which implies that $Z(\beta)\approx 1$. Consequently, for the remaining N1N-1 states, the probabilities can be approximated as: